What is the limiting reactant when 12.8 g of
CuO are exposed to 2.46 g of H2 according to
the equation
CuO(s) + H2(g) −→ Cu(s) + H2O(g) ?
1
Expert's answer
2020-04-27T01:44:56-0400

When reacting with 1 mol of copper oxide, 1 mol of hydrogen reacts


n(CuO)=m/M=12,8g/79,5g/mol=0,16mol

n(H2)=2,46g/2g/mol=1,23mol


Copper oxide is in short supply, therefore, it is a limiting agent.